Calculate the normal force, N acting in a elevator if his mass is 72.1kg and the acceleration due to gravity is 9.81m/s2. the elevator has constant velocity
Answer #1

As the velocity of elevator is constant, the acceleration is zero.

Normal force = mg = (72.1*g) = 72.1*9.8 = 706.58 N.